# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v5

@@ 4.2 Blanket is “in” the general clone
$(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})$, $n$, $\mathbb{Y}=\{y_{1},y_{2},\dots,y_{|\mathbb{Y}|}\}$
$H\leftarrow\{\}$triangleright $H$ is a histogram
@@
     $j\leftarrow\text{index}(y)\text{ such that }y_{j}=y$
     $H[j]\leftarrow H[j]+1$triangleright Map $y=y_{j}$ in Blanket to index $j$ in GC
@@
$H[|\mathbb{Y}|+1]\leftarrow n-1-|\boldsymbol{b}|$triangleright Map $y=\bot$ in Blanket to index $|\mathbb{Y}|+1$ in GC
@@ 4.3 General clone is not stronger than blanket
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X_{0})\parallel\mathcal{M}_{S}(X_{1}))\leq D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}}\parallel P_{1}^{\mathsf{B}})\leq D_{e^{\varepsilon}}(P_{0}^{\mathsf{GC}}\parallel P_{1}^{\mathsf{GC}}).$
@@
Let $f_{3}$ be the post-processing function defined in Algorithm 5. Its behavior is as follows: it processes each input sequentially. When the input is $j\in[1,k]$, it outputs a sample from $Q_{1}^{j}$; when $j=k+1$, it samples from $Q^{\text{com}}$ with probability $\mu/\beta$.
We now argue that $f_{3}(P_{0}^{\mathsf{GC}})=P_{0}^{\mathsf{B}}$. By definition, $P_{0}^{\mathsf{B}}$ outputs one sample from $\mathcal{R}(x_{1}^{0})$, and $|\boldsymbol{b}|\sim\text{Bin}(n-1,\gamma)$ samples from $\omega$. From Definition 8, $P_{0}^{\mathsf{GC}}$ draws one sample from $A_{0}$, and $n-1$ samples from $A_{2}$. Since $\text{supp}(A_{0})\subseteq\{1,\dots,k\}$, its output is always retained. The transformed output has distribution:
@@
$=\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)/\gamma+\frac{\mu}{\gamma}\cdot Q^{\text{com}}(y)$
$=\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)/\gamma+\frac{\mu}{\gamma}\cdot\frac{\gamma\omega(y)-\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)}{\mu}$
@@
for $j=1,\dots,k+1$ do
     for $count=1,2,\dots,n_{j}$ do
         if $j=k+1$ then
              $r\leftarrow\text{Bern}(\frac{\gamma-\sum_{m=1}^{k}c_{m}}{\beta})$
@@
                  $s\leftarrow_{Q^{\text{com}}}\mathbb{Y}$triangleright Sample from $Q^{\text{com}}$
@@
              $s\leftarrow_{Q^{j}_{1}}\mathbb{Y}$triangleright Sample from $Q^{j}_{1}$
@@
return $(|Y|-1,Y)$
As a result, although both $D_{e^{\varepsilon}}(P_{0}^{\mathsf{GC}}\parallel P_{1}^{\mathsf{GC}})$ and $D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}}\parallel P_{1}^{\mathsf{B}})$ serve as upper bounds on the privacy amplification in the shuffle model, the bound provided by the privacy blanket is always at least as tight as that of the general clone paradigm under any decomposition.
@@ 4.4 Simplified optimal decomposition
$S_{r_{0},r_{1}}=\left\{y\ \middle|\ \frac{\mathcal{R}(x_{1}^{0})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{0},\quad\frac{\mathcal{R}(x_{1}^{1})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{1}\right\}.$
@@
$(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})$, $n$
$H\leftarrow\{\}$triangleright $H$ is a histogram
@@
$H[\bot]\leftarrow n-1-|\boldsymbol{b}|$triangleright The residual part of GC
@@
[…]
@@
Its simplified optimal decomposition is as follows:
$\mathcal{R}(x_{1}^{0})$ $=e^{\varepsilon_{0}}p\mathbbm{1}_{x_{1}^{0}}+p\mathbbm{1}_{x_{1}^{1}}+qU,$
$\mathcal{R}(x_{1}^{1})$ $=p\mathbbm{1}_{x_{1}^{0}}+e^{\varepsilon_{0}}p\mathbbm{1}_{x_{1}^{1}}+qU,$
$\forall i\in[2,n]:\mathcal{R}(x_{i})$ $=p\mathbbm{1}_{x_{1}^{0}}+p\mathbbm{1}_{x_{1}^{1}}+qU+(e^{\varepsilon_{0}}-1)p\mathbbm{1}_{x_{i}}.$
where $p=\frac{1}{e^{\varepsilon_{0}}+k-1}$, $q=(k-2)p$, and $U$ is the uniform distribution over $[k]-\{x_{1}^{0},x_{1}^{1}\}$. This decomposition is also considered in the corrected version of the stronger clone [FMT23b].
The optimal decompositions for Binary Local Hash [WBLJ17], RAPPOR [EPK14], Optimized Unary Encoding (OUE) [WBLJ17] and Hadamard Response [ASZ19] are provided in Appendix A. It is worth noting that the optimal decompositions of BLH, RAPPOR, and OUE depend on the domain size $D$. However, as $D\to\infty$, each decomposition converges to a limiting form that is independent of $D$, with a convergence rate of $O(2^{-D})$. We provide both the exact, $D$-dependent solutions in Table 2 and the asymptotic, $D$-independent solutions in Table 3.
@@ 4.5 Optimal decompositions of the joint composition
×$\mathcal{R}^{\times}(x_{1}^{1},x_{1}^{2})$ $=\sum_{y^{1},y^{2}}a_{y^{1},y^{2}}^{\prime}\mathbbm{1}_{(y^{1},y^{2})}$
×$\mathcal{R}^{\times}(\bar{x}_{1}^{1},\bar{x}_{1}^{2})$ $=\sum_{y^{1},y^{2}}b_{y^{1},y^{2}}^{\prime}\mathbbm{1}_{(y^{1},y^{2})}$
@@
$a_{y^{1},y^{2}}^{\prime}$ =×$=\ \mathcal{R}^{\times}(x_{1}^{1},x_{1}^{2})(y^{1},y^{2})$ $\quad$=\ = $\ \mathcal{R}_{1}(x_{1}^{1})(y^{1})\cdot\mathcal{R}_{2}(x_{1}^{2})(y^{2})$ $\quad$=\ = $\ a_{y^{1}}\cdot d_{y^{2}}$
$b_{y^{1},y^{2}}^{\prime}$ =×$=\ \mathcal{R}^{\times}(\bar{x}_{1}^{1},\bar{x}_{1}^{2})(y^{1},y^{2})$ $\quad$=\ = $\ \mathcal{R}_{1}(\bar{x}_{1}^{1})(y^{1})\cdot\mathcal{R}_{2}(\bar{x}_{1}^{2})(y^{2})$ $\quad$=\ = $\ b_{y^{1}}\cdot e_{y^{2}}$
$c_{y^{1},y^{2}}^{\prime}$ $=\ \inf_{x^{1},x^{2}}\mathcal{R}^{\times}(x^{1},x^{2})(y^{1},y^{2})$ $\quad$=\ = $\ \inf_{x^{1}}\mathcal{R}_{1}(x^{1})(y^{1})\cdot\inf_{x^{2}}\mathcal{R}_{2}(x^{2})(y^{2})$ $\quad$=\ = $\ c_{y^{1}}\cdot f_{y^{2}}.$
@@
Next, we show that the Cartesian product of the simplified Optimal Decompositions of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ also forms an Optimal Decomposition of $\mathcal{R}^{\times}$. Recall that a simplified Optimal Decomposition (given by Theorem 7) is formed by grouping according to the following sets:
$S_{p_{0},p_{1}}^{1}$ =$=\left\{y\in\mathbb{Y}^{1}\,\middle|\,\frac{\mathcal{R}_{1}(x_{1}^{1})(y)}{\inf_{x}\mathcal{R}_{1}(x)(y)}=p_{0},\ \frac{\mathcal{R}_{1}(\bar{x}_{1}^{1})(y)}{\inf_{x}\mathcal{R}_{1}(x)(y)}=p_{1}\right\}$
$S_{q_{0},q_{1}}^{2}$ =$=\left\{y\in\mathbb{Y}^{2}\,\middle|\,\frac{\mathcal{R}_{2}(x_{1}^{2})(y)}{\inf_{x}\mathcal{R}_{2}(x)(y)}=q_{0},\ \frac{\mathcal{R}_{2}(\bar{x}_{1}^{2})(y)}{\inf_{x}\mathcal{R}_{2}(x)(y)}=q_{1}\right\}$
Each component in the Cartesian product of the simplified Optimal Decompositions of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ corresponds to the Cartesian product set $S_{p_{0},p_{1}}^{1}\times S_{q_{0},q_{1}}^{2}$. According to the rule of simplified Optimal Decomposition, all elements with the same ratio
@@
There is a minor gap in the proof: in the simplified optimal decomposition (Theorem 7), all components of the primary optimal decomposition that share the same ratio $a_{j}:b_{j}:c_{j}$ are merged into a single component, whereas the decomposition obtained via the Cartesian product yields a finer partition.
@@
As an example, consider the joint composition of two $k$-RR mechanisms acting on $[k]\times[k]$. The optimal decomposition is given below:
Example 4.2 (Optimal Decomposition of Two-Joint $k$-RR).
@@
where $p=\frac{1}{e^{\varepsilon_{0}/2}+k-1}$, $q=\frac{k-2}{e^{\varepsilon_{0}/2}+k-1}$, and $U$, $V$ are the uniform distributions over $[k]\setminus\{a_{1}^{0},a_{1}^{1}\}$ and $[k]\setminus\{b_{1}^{0},b_{1}^{1}\}$, respectively.
@@ 4.6 Optimal decompositions of the parallel composition
An optimal decomposition of [$\mathcal{R}^{\parallel,\boldsymbol{p}}_{[1:m]}$ is a weighted combination of the primary (or simplified) optimal decompositions of the individual local randomizers $\mathcal{R}_{i}$, with weights given by $p_{i}$.
@@
$\mathcal{R}^{\parallel,\boldsymbol{p}}(x_{1}^{0})$ $=p_{1}\sum_{y^{1}\in\mathbb{Y}^{1}}a_{y^{1}}\mathbbm{1}_{y^{1}}+p_{2}\sum_{y^{2}\in\mathbb{Y}^{2}}d_{y^{2}}\mathbbm{1}_{y^{2}},$
$\mathcal{R}^{\parallel,\boldsymbol{p}}(x_{1}^{1})$ $=p_{1}\sum_{y^{1}\in\mathbb{Y}^{1}}b_{y^{1}}\mathbbm{1}_{y^{1}}+p_{2}\sum_{y^{2}\in\mathbb{Y}^{2}}e_{y^{2}}\mathbbm{1}_{y^{2}},$
$\forall i\in[2,n],\ \mathcal{R}^{\parallel,\boldsymbol{p}}(x_{i})$ $=p_{1}\left(\sum_{y^{1}\in\mathbb{Y}^{1}}c_{j}\mathbbm{1}_{y^{1}}+\beta_{1}Q_{i}\right)+p_{2}\left(\sum_{y^{2}\in\mathbb{Y}^{2}}f_{y^{2}}\mathbbm{1}_{y^{2}}+\beta_{2}T_{i}\right).$ (9)
@@
$a_{1,y^{1}}^{\prime}$ $=p_{1}\cdot\mathcal{R}_{1}(x_{1}^{0})(y^{1}),$ $a_{2,y^{2}}^{\prime}$ $=p_{2}\cdot\mathcal{R}_{2}(x_{1}^{0})(y^{2}),$
$b_{1,y^{1}}^{\prime}$ $=p_{1}\cdot\mathcal{R}_{1}(x_{1}^{1})(y^{1}),$ $b_{2,y^{2}}^{\prime}$ $=p_{2}\cdot\mathcal{R}_{2}(x_{1}^{1})(y^{2}),$
$c_{1,y^{1}}^{\prime}$ $=p_{1}\cdot\inf_{x\in\mathbb{X}}\mathcal{R}_{1}(x)(y^{1}),$ $c_{2,y^{2}}^{\prime}$ $=p_{2}\cdot\inf_{x\in\mathbb{X}}\mathcal{R}_{2}(x)(y^{2}).$
@@
For all $\text{For all }y\in S_{r_{0}^{(i)},r_{1}^{(i)}},\quad a_{i,y^{\prime}}:b_{i,y^{\prime}}:c_{i,y^{\prime}}=r_{0}^{(i)}:r_{1}^{(i)}:1.$
@@
A classical example of parallel composition is computing the mean of an $m$-dimensional numerical vector, where $\vec{x}=(x^{1},x^{2},\dots,x^{m})\in[0,1]^{m}$. Define $R_{i}(\vec{x})=\big{(}i,x^{i}+\mathrm{Lap}(1/\varepsilon_{0})\big{)}$ for $i=1,2,\dots,m$, each satisfying $\varepsilon_{0}$-LDP. When the probability vector is uniform, i.e., $\boldsymbol{p}=(1/m,1/m,\dots,1/m)$, the mechanism [$\mathcal{R}^{\parallel,\boldsymbol{p}}_{[1,m]}$ corresponds to the setting where each user randomly selects one coordinate and reports it using the Laplace mechanism.
@@
An important theoretical advantage of parallel composition is that it subsumes Poisson subsampling, as discussed in Section 1.1. In the experimental section, we present the results of our method in computing upper and lower bounds on the privacy amplification of the $k$-RR mechanism under Poisson subsampling followed by shuffling.

### v6

@@ 4.2 Blanket is “in” the general clone
  Meta parameter: $n$, $\mathbb{Y}=\{y_{1},y_{2},\dots,y_{|\mathbb{Y}|}\}$
$(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})$
$\leftarrow\{\} \triangleright HH$ is a histogram
@@
[…]
  $H[j]\leftarrow H[j]+1$ $\triangleright$ Map $y=y_{j}$ in Blanket to index $j$ in GC
@@
$H[|\mathbb{Y}|+1]\leftarrow n-1-|\boldsymbol{b}|$ $\triangleright$ Map $y=\bot$ in Blanket to index $|\mathbb{Y}|+1$ in GC
@@ 4.3 General clone is not stronger than blanket
$D_{e^{\varepsilon}}(\mathcal{M}_{S}(X_{0}),\mathcal{M}_{S}(X_{1}))\leq D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})\leq D_{e^{\varepsilon}}(P_{0}^{\mathsf{GC}},P_{1}^{\mathsf{GC}}).$
@@
Let $f_{3}$ be the post-processing function defined in Algorithm 5. Its behavior is as follows: it processes each input sequentially. When the input is $j\in[1,k]$, it outputs a sample from $Q_{1}^{j}$; when $j=k+1$, it samples from $Q^{\text{com}}$ with a certain probability.
We now argue that $f_{3}(P_{0}^{\mathsf{GC}})=P_{0}^{\mathsf{B}}$. By definition, $P_{0}^{\mathsf{B}}$ outputs one sample from $\mathcal{R}(x_{1}^{0})$, and $|\boldsymbol{b}|\sim\text{Bin}(n-1,\gamma)$ samples from $\omega$. From Definition 8, $P_{0}^{\mathsf{GC}}$ draws one sample from $A_{0}$, and n1n-1 samples from $A_{2}$. Since $\text{supp}(A_{0})\subseteq\{1,\dots,k\}$, its output is always retained. The transformed output has distribution:
@@
$=\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)/\gamma+\frac{\gamma-\sum_{j=1}^{k}c_{j}}{\gamma}\cdot Q^{\text{com}}(y)$
$=\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)/\gamma+\frac{\gamma-\sum_{j=1}^{k}c_{j}}{\gamma}\cdot\frac{\gamma\omega(y)-\sum_{j=1}^{k}c_{j}Q_{1}^{j}(y)}{\gamma-\sum_{j=1}^{k}c_{j}}$
@@
for $i=1,\dots,k+1$ do
  for $count=1,2,\dots,n_{i}$ do
   if $i=k+1$ then
     $r\leftarrow\text{Bern}(\frac{\gamma-\sum_{j=1}^{k}c_{j}}{\beta})$
@@
      $s\leftarrow_{Q^{\text{com}}}\mathbb{Y}$ $\triangleright$ Sample from $Q^{\text{com}}$
@@
     $s\leftarrow_{Q^{i}_{1}}\mathbb{Y}$ $\triangleright$ Sample from $Q^{i}_{1}$
@@
return (|Y|1,Y)(|Y|-1,Y)
As a result, although both $D_{e^{\varepsilon}}(P_{0}^{\mathsf{GC}}||P_{1}^{\mathsf{GC}})$ and $D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}}||P_{1}^{\mathsf{B}})$ serve as upper bounds on the privacy amplification in the shuffle model, the bound provided by the privacy blanket is always at least as tight as that of the general clone paradigm under any decomposition.
@@ 4.4 Simplified optimal decomposition
$S_{r_{0},r_{1}}=\{y|\frac{\mathcal{R}(x_{1}^{0})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{0},\frac{\mathcal{R}(x_{1}^{1})(y)}{\inf_{x}\mathcal{R}(x)(y)}=r_{1}\}.$
@@
  Meta parameter: $n$
$(|\boldsymbol{b}|,Y_{|\boldsymbol{b}|})$
$\leftarrow\{\} \triangleright HH$ is a histogram
@@
$H[\bot]\leftarrow n-1-|\boldsymbol{b}|$ $\triangleright$ The residual part of GC
@@
$\forall x,x^{\prime},\ \forall y:\quad\frac{\mathcal{R}(x)(y)}{\mathcal{R}(x^{\prime})(y)}\in\{e^{-\varepsilon_{0}},\ 1,\ e^{\varepsilon_{0}}\}.$
@@
Its optimal decomposition is as follows:
$R(x_{1}^{0})$ $=e^{\varepsilon_{0}}p\mathbbm{1}_{x_{1}^{0}}+p\mathbbm{1}_{x_{1}^{1}}+qU,$
$R(x_{1}^{1})$ $=p\mathbbm{1}_{x_{1}^{0}}+e^{\varepsilon_{0}}p\mathbbm{1}_{x_{1}^{1}}+qU,$
$\forall i\in[2,n]:R(x_{i})$ $=p\mathbbm{1}_{x_{1}^{0}}+p\mathbbm{1}_{x_{1}^{1}}+qU+(e^{\varepsilon_{0}}-1)p\mathbbm{1}_{x_{i}}.$
where $p=\frac{1}{e^{\varepsilon_{0}}+k-1}$, q=(k2)pq=(k-2)p, and $U$ is the uniform distribution over $k]-\{x_{1}^{0},x_{1}^{1}\}$. This decomposition is also considered in the corrected version of the stronger clone [feldman2023arxiv].
The optimal decompositions for Binary Local Hash [3], RAPPOR [6], Optimized Unary Encoding (OUE) [3] and Hadamard Response [HR] are provided in Appendix A. It is worth noting that the optimal decompositions of these randomizers depend on the domain size $D$. However, as $D\to\infty$, each decomposition converges to a limiting form that is independent of $D$, with a convergence rate of $O(2^{-D})$. We provide both the exact, $D$-dependent solutions in Table 2 and the asymptotic, $D$-independent solutions in Table 3.
@@ 4.5 Optimal decompositions of the joint composition
$\mathcal{R}^{\times}(x_{1}^{1},x_{1}^{2})$ $=\sum_{y^{1},y^{2}}a_{y^{1},y^{2}}^{\prime}\mathbbm{1}_{(y^{1},y^{2})}$
$\mathcal{R}^{\times}(\bar{x}_{1}^{1},\bar{x}_{1}^{2})$ $=\sum_{y^{1},y^{2}}b_{y^{1},y^{2}}^{\prime}\mathbbm{1}_{(y^{1},y^{2})}$
@@
[…]
@@
Next, we show that the Cartesian product of the simplified Optimal Decompositions of $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ also forms an Optimal Decomposition of $\mathcal{R}^{\times}$. Recall that a simplified Optimal Decomposition is formed by grouping according to the following sets:
$S_{p_{0},p_{1}}^{1}$ $=\left\{y\in\mathbb{Y}^{1}\,\middle|\,\frac{\mathcal{R}_{1}(x_{1}^{1})(y)}{\inf_{x}\mathcal{R}_{1}(x)(y)}=p_{0},\ \frac{\mathcal{R}_{1}(\bar{x}_{1}^{1})(y)}{\inf_{x}\mathcal{R}_{1}(x)(y)}=p_{1}\right\}$
$S_{q_{0},q_{1}}^{2}$ $=\left\{y\in\mathbb{Y}^{2}\,\middle|\,\frac{\mathcal{R}_{2}(x_{1}^{2})(y)}{\inf_{x}\mathcal{R}_{2}(x)(y)}=q_{0},\ \frac{\mathcal{R}_{2}(\bar{x}_{1}^{2})(y)}{\inf_{x}\mathcal{R}_{2}(x)(y)}=q_{1}\right\}$
Each component in the product decomposition corresponds to the Cartesian product set $S_{p_{0},p_{1}}^{1}\times S_{q_{0},q_{1}}^{2}$. According to the rule of simplified Optimal Decomposition, all elements with the same ratio
@@
There is a minor gap in the proof: in the simplified optimal decomposition, all components of the primary optimal decomposition that share the same ratio $a_{j}:b_{j}:c_{j}$ are merged into a single component, whereas the decomposition obtained via the Cartesian product yields a finer partition.
@@
As an example, consider the joint composition of two $k$-$R$ mechanisms acting on $k]\times[k]$. The optimal decomposition is given below:
Example 4.2 (Optimal Decomposition of Two-Joint $k$-$R$).
@@
where $p=\frac{1}{e^{\varepsilon_{0}/2}+k-1}$, $q=\frac{k-2}{e^{\varepsilon_{0}/2}+k-1}$, and $U$, $V$ are the uniform distributions over $k]\setminus\{a_{1}^{0},a_{1}^{1}\}$ and $k]\setminus\{b_{1}^{0},b_{1}^{1}\}$, respectively.
@@ 4.6 Optimal decompositions of the parallel composition
An optimal decomposition of $\mathcal{R}^{\parallel,\boldsymbol{p}}_{[1:m]}$ is a weighted combination of the primary (or simplified) optimal decompositions of the individual local randomizers $\mathcal{R}_{i}$, with weights given by $p_{i}$.
@@
$\mathcal{R}^{\parallel}(x_{1}^{0})$ $=p_{1}\sum_{y^{1}\in\mathbb{Y}^{1}}a_{y^{1}}\mathbbm{1}_{y^{1}}+p_{2}\sum_{y^{2}\in\mathbb{Y}^{2}}d_{y^{2}}\mathbbm{1}_{y^{2}},$
$\mathcal{R}^{\parallel}(x_{1}^{1})$ $=p_{1}\sum_{y^{1}\in\mathbb{Y}^{1}}b_{y^{1}}\mathbbm{1}_{y^{1}}+p_{2}\sum_{y^{2}\in\mathbb{Y}^{2}}e_{y^{2}}\mathbbm{1}_{y^{2}},$
$\forall i\in[2,n],\ \mathcal{R}^{\parallel}(x_{i})$ $=p_{1}\left(\sum_{y^{1}\in\mathbb{Y}^{1}}c_{j}\mathbbm{1}_{y^{1}}+\beta_{1}Q_{i}\right)+p_{2}\left(\sum_{y^{2}\in\mathbb{Y}^{2}}f_{y^{2}}\mathbbm{1}_{y^{2}}+\beta_{2}T_{i}\right).$ (9)
@@
$a_{1,y^{1}}^{\prime}$ $=p_{1}\mathcal{R}_{1}(x_{1}^{0})(y^{1}),$ $a_{2,y^{2}}^{\prime}$ $=p_{2}\mathcal{R}_{2}(x_{1}^{0})(y^{2}),$
$b_{1,y^{1}}^{\prime}$ $=p_{1}\mathcal{R}_{1}(x_{1}^{1})(y^{1}),$ $a_{2,y^{2}}^{\prime}$ $=p_{2}\mathcal{R}_{2}(x_{1}^{1})(y^{2}),$
$c_{1,y^{1}}^{\prime}$ $=p_{1}\inf_{x\in\mathbb{X}}\mathcal{R}_{1}(x)(y^{1}),$ $c_{2,y^{2}}^{\prime}$ $=p_{2}\inf_{x\in\mathbb{X}}\mathcal{R}_{2}(x)(y^{2}).$
@@
$\text{For all }y\in S_{r_{0}^{(i)},r_{1}^{(i)}},\quad a_{i,y^{\prime}}:b_{i,y^{\prime}}:c_{i,y^{\prime}}=r_{0}^{(i)}:r_{1}^{(i)}:1.$
@@
A classical example of parallel composition is computing the mean of an $m$-dimensional numerical vector, where $\vec{x}=(x^{1},x^{2},\dots,x^{m})\in[0,1]^{m}$. Define $R_{i}(\vec{x})=\big(i,x^{i}+\mathrm{Lap}(1/\varepsilon_{0})\big)$ for $i=1,2,\dots,m$, each satisfying $\varepsilon_{0}$-LDP. When the probability vector is uniform, i.e., $\boldsymbol{p}=(1/m,1/m,\dots,1/m)$, the mechanism $\mathcal{R}^{\parallel,\boldsymbol{p}}_{[1,m]}$ corresponds to the setting where each user randomly selects one coordinate and reports it using the Laplace mechanism.
@@
An important theoretical advantage of parallel composition is that it subsumes Poisson subsampling, as discussed in Section 1.1. In the experimental section, we present the results of our method in computing upper bounds on the privacy amplification of the $k$-$R$ mechanism under Poisson subsampling followed by shuffling.